OpenStudy (anonymous):

What is the value of w? 5(9 − 3) = w • 9 – 5 • 3 w =

OpenStudy (anonymous):

remember PEMDAS. first distribute the 5 45-15=w*9-5*3 then multiple w and 9, and 5 and 3 45-15=9w-15 combine like terms 30=9w-15 start getting 9 by itself 15=9w 15/9 = 5/3 = w